OpenStudy (anonymous):

A coffee can has a volume of 137.375 cubic inches and a height of 7 inches. Find the circumference of this coffee can? the answers are A. 15.7 in B. 6.25 in C. 12.5 in D. 19.625in

OpenStudy (anonymous):

can you help me

OpenStudy (souvik):

coffee can is a cylinder... its volume V=pi*r^2*h here r =radius of the base and h=height find r first

OpenStudy (anonymous):

how would i do that

OpenStudy (souvik):

do u have a calculator?

OpenStudy (anonymous):

yes

OpenStudy (souvik):

then use it

OpenStudy (anonymous):

i got 19.625 is that right

OpenStudy (souvik):

137.375=3.14*r^2*7 r=?

OpenStudy (anonymous):

19.625in

OpenStudy (souvik):

i got 2.5 in sorry there is a disturb in network

OpenStudy (anonymous):

thats not oone of the answers

OpenStudy (anonymous):

but when i divide 137.375 in to 7 i get 19.625in

OpenStudy (souvik):

that is not your answer.. the circumference of this coffee can is 2*pi*r

OpenStudy (anonymous):

look in the questions you will see the 4 answers i have to choose from

OpenStudy (souvik):

i got r=2.5 in so circumference=2*3.14*2.5=15.7 in

OpenStudy (anonymous):

15.7 is the answer if you go up to the question the answer you got would be A

OpenStudy (souvik):

yes..

OpenStudy (anonymous):

thanck you for your help

OpenStudy (souvik):

welcome